OpenStudy (anonymous):

Create a quadratic polynomial function f(x) and a linear binomial in the form (x − a). Part 1. Show all work using long division to divide your polynomial by the binomial. Part 2. Show all work to evaluate f(a) using the function you created. Part 3. Use complete sentences to explain how the remainder theorem is used to determine whether your linear binomial is a factor of your polynomial function

OpenStudy (anonymous):

@dmndlife24 @SolomonZelman

OpenStudy (anonymous):

@crissyyyxoxo

OpenStudy (misssunshinexxoxo):

f(x) = x^3 - 3x^2 - x + 3., and a linear function g(x) = (x - 1). Part 1. Divide f(x) by g(x). You must do it yourself using the long division you have learned f(x)/g(x) = (x^3 - 3x^2 + x + 3)/(x - 1) = x^2 - 2x + 3 (no remainder) Part 2: a = 1 -> f(a) = f(1) = 0

OpenStudy (misssunshinexxoxo):

The remainder theorem is based on synthetic division, which is the process of dividing a polynomial f(x) by a polynomial D(x) and finding the remainder. This is written as , where f(x) is the dividend, Q(x) is the quotient, D(x) is the divisor, and R(x) is the remainder.

OpenStudy (anonymous):

^^ ahh @misssunshinexxoxo caught it b4 me lol

OpenStudy (misssunshinexxoxo):

@Mathania if you got questions please let me know